CAT Chat Minutes

June 12, 1998

Information/Follow-up Items

1. A fire alarm sounded on Thursday, June 11. While the alarm sounded inside the tunnel, we could hear it in some of the sectors on the experiment hall floor. In the capacity of "emergency area monitor" we moved all personnel to LOM doorway. It would be good to know the required actions, should this recur.

Generally, CAT-selected CAT personnel are given the responsibility for "emergency area monitor" for a specific LOM. They are not required to act if the alarm is sounding either in the storage ring tunnel or on the experiment hall floor. The "emergency area monitors" for the experiment hall floor are usually the Floor Coordinators. They will give instructions in such circumstances. Normally this will require personnel to move to the LOM. If the alarm is sounded in the LOM, everyone in the LOM should vacate the LOM and move to the outside of the building. Please note that the alarms are normally followed by voice alerts that provide instructions to follow.

APS Operational Status Update

The availability overall for the run stands at 93%; the past week has been at slightly over 93%. At the end of last week, there was a water leak on the rf load, and yesterday there were false fire alarms in the storage ring due to a faulty detector. A trip on rf-1 this morning has resulted in a switch to rfs 2 and 4. The server "oxygen" was down on Wednesday of this week due to a hardware problem.

Monday, June 15 is an ASD studies day. The beam will be used on Tuesday, June 16 from 8:00 a.m. for shielding verification. We expect the user beam by about 2:00 p.m.

Information

The network meetings (covering network configuration, upgrades, and connectivity issues) will soon be held on a regular basis.

The computer support pages discussed last week have been made accessible to users (they were originally accessible to APS personnel).

Sector 33 has reported seeing a roughly one-half Hz beam size fluctuation in the vertical plane. The fluctuation is not totally periodic and has been seen at multiple gap settings. No correlation to any storage ring parameters has been found. CATs are asked to report any similar observations to the APS.

It was announced that SLAC suffered a serious break-in to their entire network. The breach may necessitate the shutting down of the entire link to the outside world for as long as a week. Details will be passed along when they become available.